X Company has two production departments, A and B. The following is budgeted information for all of its products in 2019, and actual information for one of them, Product X:

Department A All Products Product X
   Overhead $7,900,000 --   
   Direct labor hours 60,000 788  
   Machine hours 120,000 1,080  
   Units produced 58,000 560  
Department B
   Overhead $1,900,000 --   
   Direct labor hours 40,000 142  
   Machine hours 130,000 900  
   Units produced 46,000 560  

Using a single-driver allocation system, with machine hours as the driver, how much overhead was allocated to Product X [round overhead rate(s) to two decimal places]?
Answer #1
Total Overhead 9800000 =7900000+1900000
Divide by Total Machine hours 250000 =120000+130000
Overhead rate 39.20
Total Machine hours Product X 1980 =1080+900
X Overhead rate 39.20
Overhead allocated to Product X 77616